OpenStudy (anonymous):

Diff the following functions, 1.) tan^-1((a+x)/(1-ax)) 2.)tan^-1((a+btanx)/(b-atanx)) My question is can you substitute for constants? like a=tan(alpha)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

no i don't think so

OpenStudy (anonymous):

oh

OpenStudy (misty1212):

i think the idea of this question is this: once you find the derivative of the first one, the derivative of the second will be easy by the chain rule

OpenStudy (anonymous):

These questions are to be done by trigonometric subtitutions but I'm just stuck on the 2nd one, and I just found I did the first one wrong because you can't sub for constants according to you

OpenStudy (misty1212):

hmm i may be wrong about this, i have never heard of a trig sub for a derivative, only for an integral a quick wolfram check tells me the derivative of the first one is identical to the derivative of \(\tan^{-1}(x)\)

OpenStudy (misty1212):

it is not at all clear to me why however if you do it the long way you get \(\frac{1}{1+x^2}\)

OpenStudy (anonymous):

Oh? Here's an example, y=sin^-1(root(1-x^2)) x=cos(theta) theta=cos^-1(x) y=sin^-1(root(1-cos^2(theta))) y=sin^-1(root(sin^2(theta))) y=sin^-1(sin(theta)) y=theta y=cos^-1(x) dy/dx= -1/root(1-x^2)

OpenStudy (misty1212):

i see

OpenStudy (anonymous):

Sometimes in inverse trig functions, you can sub and then derivative, it's en exercise in my textbook

OpenStudy (dumbcow):

heres a reference of derivatives for inverse trig. https://en.wikipedia.org/wiki/Inverse_trigonometric_functions#Derivatives_of_inverse_trigonometric_functions then use chain rule

OpenStudy (misty1212):

@dumbcow this is somehow different

OpenStudy (misty1212):

i mean you can do it they way you suggested of course, but there is some trick involved using a trig sub

OpenStudy (anonymous):

That is good and all Professor, but these questions are to be done with sub, doing the direct way would take time

OpenStudy (dumbcow):

oh i see... i think the trig sub just allows you to derive the formula for derivative

OpenStudy (misty1212):

so would the sub be \(x=\tan(\theta)\)?

OpenStudy (anonymous):

Yeah there is probably some trick, but first things first I would like to know if you could sub for constants or not, because otherwise Im just going down the wrong road

OpenStudy (misty1212):

no i don't think so, i think you sub for the variable

OpenStudy (misty1212):

let me try with pencil and paper and see if something nice comes out of it

OpenStudy (dumbcow):

what is alpha?

OpenStudy (anonymous):

The first question i did like this a=tan(alpha), x=tan(beta) alpha=tan^-1(a), beta=tan^-1(x) y=tan^-1(tan(alpha)+tan(beta)/1-tan(alpha)tan(beta)) y=tan^-1(tan(alpha+beta)) y=alpha+beta y=tan^-1(a)+tan^-1(x) dy/dx=1/1+x^2

OpenStudy (anonymous):

alpha is variable of substitution I guess, doesn't really matter as we would sub back alpha for x

OpenStudy (misty1212):

i think you best leave \(a\) alone as it is a constant, but i could be wrong

OpenStudy (dumbcow):

hmm i think that works ... http://www.wolframalpha.com/input/?i=+differentiate+arctan%28%28a%2Bx%29%2F%281-ax%29%29 yep:)

OpenStudy (anonymous):

Yeah I was confused about that, but you can see how easily a solution comes out if you can sub for constants...

OpenStudy (misty1212):

your answer is definitely right

OpenStudy (misty1212):

yeah i guess you can the constants drop out

OpenStudy (anonymous):

Yeah, I checked the answer. It IS right but I don't know if the method IS right

OpenStudy (dumbcow):

@misty1212 , rewriting a constant in different form is always ok k = e^C kinda thing

OpenStudy (misty1212):

so trick was recognizing the addition angel formula for tangent

OpenStudy (anonymous):

yeah, it looks like the tan(2x) identity if you notice carefully

OpenStudy (dumbcow):

nice job, i didn't notice that at all

OpenStudy (misty1212):

OH! light goes on learn something new every day

OpenStudy (anonymous):

sorry not tan(2x), tan(x+y)

OpenStudy (misty1212):

yeah got that wow

OpenStudy (misty1212):

so for the second one, use your answer to first one plus chain rule right? or is there some other thing going on?

OpenStudy (anonymous):

you probably factor out a or b from both numberator and denominator then sub

OpenStudy (anonymous):

something like 1+(b/a)tanx, something like that

OpenStudy (dumbcow):

sorry i was away, yes make that sub ---> b/a = tan(alpha) use same identity as last one then use chain rule when doing derivative answer is 1

OpenStudy (anonymous):

@misty1212 @dumbcow Yeah I solved it, took out b from both numerator and denominator \[y=\tan^-1(\frac{ b(\frac{ a }{ b }+tanx }{ b(1-\frac{ atanx }{ b }) }) \] after that \[ \frac{ a }{ b }=\tan \alpha\] \[\alpha=\tan^-1\frac{ a }{ b }\] \[y=\tan^-1(\frac{ \tan \alpha+tanx }{ 1-\tan \alpha tanx })\] \[y=\tan^-1(\tan (\alpha+x))\] \[y=\alpha+x\]\[y=\tan^-1\frac{ a }{ b }+x\]\[\frac{ dy }{ dx }=1\]

OpenStudy (dumbcow):

:)